OpenStudy (anonymous):

Find the equation, in standard form, of the line passing through the points (3,-4) and (5,1).

OpenStudy (moazzam07):

First find the slope using equation %28y2+-+y1%29%2F%28x2+-+x1%29 Make (5,1) = (x1,y1) (-2,3) = (x2,y2) Substituting %28y2+-+y1%29%2F%28x2+-+x1%29 = %283+-+1%29%2F%28-2+-+5%29=%282%29%2F%28-7%29 Now use y = mx + b as a template. m = %282%29%2F%28-7%29 as just determined. for y and x, use the point (5,1). x=5 and y =1. Substitute these values in y = mx + b y = mx + b ; 1 = (-2/7)5 + b 1 = -10/7 + b add 10/7 to each side 1 + 10/7 = b 17/7 = b So our equation is y = (-2/7)x + 17/7

OpenStudy (moazzam07):

you can do it the same way but with your problem

OpenStudy (anonymous):

\[5 x-2 y=23 \]"Find the equation, in standard form . . ."